Suppose we have a production function Q = 16.7K - K^2. If marginal revenue is $10, what is the marginal revenue product of capital when K=4.

Hint: Round your answer to two decimal places.
Answer #1

Answer

marginal produc is change in total product and found by differentiating production function

MPK=dQ/dK=16.7-2K

K=4

MPK=16.7-2*4

MPK=8.7

MRPK=MPK*MR

MRPK=8.7*4

MRPK=34.8

the marginal revenue product of capital when K=4 is $34.80